Fast ALS was still a problem tonight. I don't think high frequency ALS noise saturating the PC drive is the issue; I put two 10k poles before the CM board (shooting for just 2-3kHz bandwidth), and the PC drive levels would be stable and low up until the lockloss, which was always conincident with a step in the AO gain.
After working with that for a few hours, we turned back to our more standard locking attempts. First, we dither aligned the PRMI, and then centered the REFL beam on REFL11. It's hard to say for certain, but we may have been a little close to the edge of the PD. The only other thing that differed from Monday's attempts was using 6dB less AO gain when trying the up the overall gain.
The script now reliably breaks through to stable high powers, we had a handful of pure-RF locks tonight. The digital DARM gain needs tuning, and the CARM bandwidth still isn't at its final state, but these are very tractable. Off the top of my head, the way forward now includes:
Unrelated: I feel that the PRC angular FF may have deteriorated a bit. I'm leaving the PRC locked on carrier to collect data for wiener filter recalculation.